# An Introduction on Solar Imaging Data Analytic Challenges and Opportunities for Statisticians

### v1

@@ 4.7 Research-to-Operation (R2O)
From SWPC/NASA, it is defined that “Research-to-Operations-to-Research (R2O2R) refers to the cyclical process by which basic research endeavors (R), having been identified as having the potential for improving forecasting capabilities, are matured, in a targeted way, toward a formal operational implementation (O) and, once operationalized, subsequent needs for refinements are conveyed back to the research community (R).” Data-drive solar eruption forecasting is motivated by and finally will serve operational use. The research-to-operation takes far more validation and testing than “a successfully trained model”. We need to handle real-time forecasting updates and model evolution over time before being able to put data-driven models into operation. This current solar cycle demonstrates quite intense Sun activities and thus can serve as a great test-bed for operational forecasting models.
@@ 4.8 Other Space Weather Phenomena
Finally, we want to mention that there are many other relevant space weather phenomena that can benefit from data-driven approaches, such as the geomagnetic index prediction (Iong et al., 2022, 2024; Ren et al., 2020) and the total electron content map reconstruction (Sun et al., 2022a, 2023a) and prediction (Sun et al., 2024; Wang et al., 2023). Camporeale (2019) gives a nice review of machine learning challenges in space weather from nowcasting and forecasting perspectives.
@@ 5 Summary
The authors hope this short introduction to solar imaging data can generate more interest in the statistics community to develop theory, methods, and computational algorithms to tackle the thorny data-analytic challenges presented by solar imaging data and more broadly space weather observations.

### v2

@@ 4.7 Research-to-Operation (R2O)
From SWPC/NASA, it is defined that “Research-to-Operations-to-Research (R2O2R) refers to the cyclical process by which basic research endeavors (R), having been identified as having the potential for improving forecasting capabilities, are matured, in a targeted way, toward a formal operational implementation (O) and, once operationalized, subsequent needs for refinements are conveyed back to the research community (R).” Data-drive solar eruption forecasting is motivated by and will finally serve operational use. The research-to-operation takes far more validation and testing than “a successfully trained model”. We need to handle real-time forecasting updates and model evolution over time before being able to put data-driven models into operation. This current solar cycle demonstrates quite intense Sun activities and thus can serve as an excellent test-bed for operational forecasting models.
@@ 4.8 Other Space Weather Phenomena
Finally, we want to mention that many other relevant space weather phenomena can benefit from data-driven approaches, such as the geomagnetic index prediction (Iong et al., 2022, 2024; Ren et al., 2020) and the total electron content map reconstruction (Sun et al., 2022a, 2023a) and prediction (Sun et al., 2024; Wang et al., 2023). Camporeale (2019) gives an excellent review of machine learning challenges in space weather from nowcasting and forecasting perspectives.
@@ 5 Summary
The authors hope this short introduction to solar imaging data can generate more interest in the statistics community to develop theory, methods, and computational algorithms to tackle the thorny data-analytic challenges presented by solar imaging data and, more broadly, space weather observations.